OpenStudy (anonymous):

What are the factored forms of the following expressions? A. 8g^2+6g-9 B. 20x^2+22x-12

OpenStudy (noorik.):

A) 8g^2+6g-9 = (4g-3)(2x+3) b) 20x^2+22x-12 = 2 (10x^2+11x-6) = 2 (5x-2)(2x+3)

OpenStudy (anonymous):

thanks man!

OpenStudy (noorik.):

you are welcome